# See, Imagine, Plan: Discovering and Hallucinating Tasks from a Single Image

### v2

@@ Abstract
Humans can not only recognize and understand the world in its current state but also envision future scenarios that extend beyond immediate perception. To resemble this profound human capacity, we introduce zero-shot task hallucination—given a single RGB image of any scene comprising unknown environments and objects, our model can identify potential tasks and imagine their execution in a vivid narrative, realized as a video. We develop a modular pipeline that progressively enhances scene decomposition, comprehension, and reconstruction, incorporating VLM for dynamic interaction and 3D motion planning for object trajectories. Our model can discover diverse tasks, with the generated task videos demonstrating realistic and compelling visual outcomes that are understandable by both machines and humans. All code, data, and additional videos are available at our project page: https://dannymcy.github.io/zeroshot_task_hallucination/
“A rock pile ceases to be a rock pile the moment a single man contemplates it, bearing within him the image of a cathedral.” —Antoine de Saint-Exupéry
Figure 1: We present a model for zero-shot task hallucination. With a single RGB image of any scene comprising unknown environments and objects, we identify potential tasks (task discovery) and imagine their execution (manipulation) in a vivid narrative, realized as a video. We also provide the trajectories of the executed tasks.
@@ 1 Introduction
The theory of mental time travel [59] highlights the human ability to recall past knowledge and anticipate future scenarios through visual perceptions. For example, the sight of a fallen chair immediately prompts in our mind an imagined rehearsal to straighten it back up. Equipping intelligent agents with this level of imaginative foresight and the ability to visualize such imagination could open up exciting possibilities, such as serving interactive tutors, offering step-by-step visual guidance on tasks, or teaching robots to autonomously discover and undertake tasks in new environments.
In this work, we aim to resemble this profound human capability to see, imagine, and plan, by utilizing the recent advancements of large foundation models [7, 47]. We introduce the problem of zero-shot task hallucination. Given a single RGB image of any scene comprising unknown environments and objects, our objective is to identify potential tasks (task discovery) and imagine their execution (manipulation) in a vivid narrative, realized as a video, as shown in Fig 1. The proposed tasks within the 3D scenes should be diverse yet reasonable. Moreover, the generated task videos should be geometric-aware and photo-realistic. The object motion should be feasible, fine-grained, and consistent to ensure the task execution is interpretable by both humans and machines.
The challenge of zero-shot task hallucination is threefold. First, the model must encompass strong capabilities of scene understanding. This includes not only the ability to recognize arbitrary objects from diverse scene settings, but also the understanding of interrelations between objects to propose reasonable tasks. Second, the model must have a 3D understanding of the scene so that the planned task execution follows the given spatial constraints. Finally, this hallucinated task execution should be presented by a human-interpretable fashion (e.g., a video). The closest existing approaches to addressing these challenges lie in video generation and 3D-aware image editing fields. Yet, no current solution seamlessly combines high-quality task execution visualization with geometric awareness. State-of-the-art (SOTA) video generation models [52, 32] are confined to 2D pixel space. They show satisfactory performance only in short sequences and struggle with the precision needed for detailed, prolonged manipulation tasks. Leading 3D-aware image editing models [18, 48] fail to maintain smooth transitions and consistency across interpolated frames. Moreover, both approaches lack the ability to understand 3D scenes and the capability to process detailed, geometry-aware prompts (e.g., covering the pot with the pot lid), which are needed for manipulation tasks.
Our solution addresses these challenges with a modular pipeline that progressively enhances scene decomposition, comprehension, and reconstruction, incorporating Vision-Language Model (VLM) for dynamic interaction and 3D motion planning for object trajectories, producing geometric-aware task videos. To understand the image scene, we use VLM to identify interactive objects and propose context-dependent tasks in a role-play manner, complemented by language-guided segmentation and repainting models to obtain occlusion-free object masks. Elevating 2D understanding to 3D, we use depth estimation and single-view 3D reconstruction models to generate a semi-reconstructed 3D scene, with the full 3D representation of foreground objects and the background as a plane. With the reconstructed 3D scene, we introduce a novel axes-constrained 3D planning approach that enables VLM to plan the motion of objects for given tasks by specifying waypoints.
Through the combination of traditional path planning algorithms, our model generates complete, feasible, and natural trajectories from merely a single image observation. With the entire framework being fully modularized, each component can be easily replaced with the latest improvements within its specific domain.
@@
    We study zero-shot task hallucination—the capability for models to discover and propose possible tasks and plans of execution given a single image.
@@
    We devise a plug-and-play framework that leverages large pretrained VLM and 3D reconstruction model, combining with traditional path planning algorithms to provide geometric-aware trajectories for diverse tasks.
@@
    We show our model can convert these task plans into human-interpretable formats such as videos for various potential applications, supported by extensive experiments.
@@ 2 Related Work
3D-Aware Image Editing   The introduction of diffusion models [24, 56] has created a plethora of high-performing aesthetic text-to-image generative models [24, 54, 3, 53, 50]. Various editing methods soon followed to bring more controls over the images generated. [6, 23, 49] enables edits to real-world images with additional words, while [19] allows fine-grained controls over the pixel space.
Several works also explored precise edits over the 3D space (e.g., object rotation, translation, illumination, non-rigid shape changes) [44, 12, 57, 48]. Zero-1-to-3 [40], ViewNETI [8], and Continuous 3D Words [12] enable rotations of an object in the image, but are generally single-object centric. 3DIT uses large-scale 3D datasets [18] to simulate and learn scene arrangements, while Diffusion Handles [48] elevates the image editing space to 3D via depth estimation.
Inspired by these works, and in light of the complicated tasks that could be discovered in our pipeline, and the recent advancements of many generalizable single-image object reconstruction techniques [39, 38, 2], we propose to explicitly reconstruct part of the scene to perform edits and rerender back to 2D.
Task Discovery   Previous research on robotic agent targeted predefined tasks [27, 42] or environments [20, 61]. Very recently, a few works [66, 1] proposed the concept of task discovery/proposal. RoboGen [66] presented the idea of generative simulation, in which they leveraged foundation models to generate robotic learning data by configuring new scenes and proposing robotic tasks in physics engines. AutoRT [1] observed and proposed tasks with a large fleet of robots. Nevertheless, both of these works made the underlying assumption that the scenes either created or observed are suitable for operation by certain robot types. Our work diverges from these constraints and is applicable to all real scenes. Moreover, our method requires only a single RGB image as observation/input.
Spatial Reasoning and Planning with VLMs   With the internet-scale training data and the birth of LLMs [7, 47], VLMs are encompassed with reasoning and scene understanding capabilities [11, 31, 37]. Several works build from the foundation and establish datasets to help with spatial reasoning/understanding [28, 36, 44]. As a concurrent work, Spatial VLM [10] enhanced standard VLMs with the ability to reason and answer questions requiring spatial understanding. In this work, we take one step further—asking VLMs to not only reason about the spatial information, but also plan the trajectory for tasks/actions in 3D space.
@@ 3 Method
Figure 2: The proposed framework. Our plug-and-play framework is fully modularized and zero-shot. Each module can be easily replaced with the latest improvements. Exact prompts for VLM are presented in Appendix. J.
Given a single RGB image $I\in\mathbb{R}^{H\times W\times 3}$ of a scene with $K$ unknown objects, our goal is to discover $Q$ tasks suitable for execution by a robotic/human hand, and generate $Q$ task videos, each visually depicting the execution of these tasks.
Our fully modularized framework progressively enhances scene decomposition, comprehension, and reconstruction, incorporating VLM for dynamic interaction and 3D motion planning for object trajectories, producing geometric-aware task videos (Fig. 2). It begins by understanding 2D image scene (Sec. 3.1). VLM identifies interactive objects and proposes context-dependent tasks through role-play, complemented by a two-step segmentation to acquire occlusion-free object masks with language-guided segmentation and repainting models. We then reconstruct the 3D scene in a semi-reconstructed fashion leveraging single-view 3D reconstruction and depth estimation models (Sec. 3.2). Our method for camera pose and object poses and sizes estimation does not require camera intrinsic matrix and uses only normalized depth, thus can be generalized to any input image. With the reconstructed 3D scene, we conduct in-context learning to let VLM understand the 3D spatial context. Subsequently, we introduce a novel axes-constrained 3D planning approach that enables VLM to plan the object motion based on the proposed tasks by specifying waypoints (Sec. 3.3). To generate complete, feasible, and natural trajectories, we use traditional robotics path planning and optimization methods. Finally, we execute the tasks within our reconstructed 3D scene following the planned paths, and render as videos.
3.1 Understanding Objects and Context of 2D Image Scene
3.1.1 Interactive Objects Identification
We start with querying VLM for intractable objects identification given $I$. We explicitly ask VLM to identify and name useful objects in term of affordances. Note that we exclude background objects (e.g., window, ceiling, stair), as they tend to be fixed and typically offer limited interaction opportunities. We also prompt VLM to provide concise descriptions, capturing the distinct color and texture of each identified object. This step is vital for differentiating between items of similar categories and appearances.

### v5

# SpatialPIN: Enhancing Spatial Reasoning Capabilities of Vision-Language Models through Prompting and Interacting 3D Priors

@@ Abstract
Current state-of-the-art spatial reasoning-enhanced VLMs are trained to excel at spatial visual question answering (VQA). However, we believe that higher-level 3D-aware tasks, such as articulating dynamic scene changes and motion planning, require a fundamental and explicit 3D understanding beyond current spatial VQA datasets. In this work, we present SpatialPIN, a framework designed to enhance the spatial reasoning capabilities of VLMs through prompting and interacting with priors from multiple 3D foundation models in a zero-shot, training-free manner. Extensive experiments demonstrate that our spatial reasoning-imbued VLM performs well on various forms of spatial VQA and can extend to help in various downstream robotics tasks such as pick and stack and trajectory planning.
@@ 1 Introduction
Equipping vision-language models (VLMs) the capacities of spatial reasoning unlocks exciting applications, such as general-purpose reward annotation [52], robotic data generation [61], and grounding 3D object affordances [26, 38]. However, the spatial reasoning capabilities of VLMs on fine-grained spatial understanding tasks are somewhat limited. Current state-of-the-art (SOTA) spatial reasoning-enhanced VLM [12] is mostly tested on spatial visual question answering (VQA), such as determining objects’ relative positions and orientations; experiments on higher-level tasks, such as scene comparisons and trajectory planning, which require more nuanced comprehension, are underexplored.
Many works enhance the spatial reasoning capabilities of VLMs by training/fine-tuning them on standard spatial VQA datasets [12]. As a result, VLMs primarily learn surface-level associations between image-text-data triplets. Given the scarcity and difficulty of obtaining spatially rich embodied data or high-quality human annotations for 3D-aware queries, we hypothesize that these VLMs may not generalize to questions outside their dataset distribution or adapt to more challenging tasks that require an advanced level of spatial understanding.
Figure 1: We present SpatialPIN, a framework to enhance the spatial reasoning capabilities of VLMs through prompting and interacting with 3D priors in a zero-shot, training-free manner.
Recent studies [73, 7, 65, 69] in image space understanding show that VLMs, equipped with internet-scale language knowledge, and multimodal foundation models capture complementary knowledge that can be combined to conduct new tasks spanning both modalities without additional training. Given the recent advancements in 3D foundation models [4, 41, 29], this work explores whether there exists an alternative approach to enhance VLMs with higher-level spatial-awareness by incorporating 3D priors from these models.
To this end, we propose SpatialPIN, a framework that utilizes progressive prompting and interactions between VLMs and 2D/3D foundation models as “free lunch” to enhance spatial reasoning capabilities in a zero-shot, training-free manner. By using these foundation models to decompose, comprehend, and reconstruct an explicit 3D representation, SpatialPIN grasps the core understanding of the 3D space presented by the 2D image. This allows generalizations to various 3D-aware tasks, from VQAs to 3D trajectory planning.
We provide an extensive empirical study combining multiple off-the-shelf and handcrafted datasets, ranging from fundamental spatial questions regarding relative positions and orientations to providing fine-grained 3D information on objects’ locations, sizes, inclinations, and dynamic changes, and plan for robotics tasks with full 3D trajectories. Results show that this straightforward approach significantly outperforms SOTA VLMs trained from extensive spatial VQAs (see SpatialPIN examples in Figure 1), consolidating our belief that a truly 3D-aware VLM can actually be imbued by simply injecting explicit, fundamental knowledge of the 3D scene. With the entire framework being fully modularized, each component can be easily replaced with the latest improvements within its specific domain.
@@
    We investigate the problem of equipping VLMs with 3D reasoning capabilities without fine-tuning on large spatial VQA datasets.
@@
    We propose SpatialPIN, a modular plug-and-play framework that progressively enhances VLM’s 3D reasoning capabilities by prompting and interacting with 3D foundational models.
@@
    We show that SpatialPIN unlocks 3D-aware applications including spatial VQA and both classic and novel robotics tasks, supported by extensive experiments.
@@ 2 Related Work
VLM Grounding   With the recent birth of powerful LLMs and VLMs [8, 40, 2], the task of VLM grounding, or combining generative language models with real-world data to adapt to specific cases, has gained significant popularity. Several recent works focused on fine-tuning these LLMs for a wide range of downstream applications, such as interactive decision making [34], multi-task agents [62], or even tasks in interactive environments [67, 11]. A close work to ours is Socratic Model [73], a framework of combining multiple foundation models to unleash LLMs in downstream tasks. However, this work still focuses on tasks in 2D pixel space understanding of images. There remains many challenges in the 3D world to combine information for full scene understanding, which we hope to tackle in our paper.
VLM Spatial Understanding   Many VLMs encompass the ability of image-space reasoning and understanding [13, 33, 40]. There are even efforts in incorporating these understandings into image space manipulations and editting [7, 65]. However, the current ability of VLMs to fully understand a 3D scene and the potential interactions within this scene is still rather limited. Several works build from this foundation and establish datasets to help with spatial reasoning/understanding [30, 39, 45]. Recently, SpatialVLM [12] proposed fine-tuning a VLM on 3D-VQA datasets to enhance the precision of VLMs on 3D understanding tasks. Nevertheless, using a 3D-VQA dataset only provides a partial picture to the complete 3D understanding of an image, and could lead to suboptimal performances under out-of-distribution tasks. In this work, we hope to introduce holistic 3D information from multiple 3D foundation models via prompting and interactions as a way to enhance VLMs with a comprehensive 3D understanding given RGB inputs.
Figure 2: SpatialPIN. Our plug-and-play framework is fully modularized and designed for zero-shot deployment. Each module can be easily replaced with the latest updates. Exact prompts for VLMs are in Appendix.
@@ 3 Method
Given an RGB image $I\in\mathbb{R}^{H\times W\times 3}$ of a scene with $K$ unknown objects and a spatial task $Q$, our goal is to inspire VLMs with spatial reasoning capabilities and solve $Q$ with fine-grained 3D understanding. To prevent the models from overfitting to the standard problems from spatial VQA datasets [12], we hope to derive a method that utilizes fundamental 3D foundation models to provide explicit scene understandings, then leverage the generalization capabilities of VLMs to tackle unforeseen tasks—all within a zero-shot, training-free manner.
Our modular pipeline, SpatialPIN, enhances VLMs’ spatial understanding of an image through progressive interactions with the scene decomposition, comprehension, and reconstruction processes with prompting. For image scene understanding (Sec. 3.1), we use VLM to describe objects by appearance and 2D location, complemented by language-guided segmentation and repainting models to obtain occlusion-free object masks. Elevating 2D understanding to coarse 3D (Sec. 3.2), we use metric depth estimation and perspective fields to estimate the 3D scene size and conduct perspective canonicalization with VLM. For fine-grained 3D understanding (Sec. 3.3), we partially reconstruct the 3D scene, with the full 3D representation of foreground objects and the background as a plane. With the reconstructed 3D scene, we summarize spatial information and prompt it to the VLM for various downstream tasks.
3.1 2D Image Scene Understanding
Prompting: Objects Understanding by Constraining   We start with querying VLM to identify and understand objects given $I$. We explicitly ask VLM to describe the objects by precise color, texture, and 2D spatial locations. This step is vital for two reasons: 1) enhance VLM’s understanding of the objects, 2) differentiate between items of similar or identical categories and appearances.
As a concrete example, given the left image of Fig. 2, VLM outputs: ‘‘
object 0: laptop of color rose gold, texture metallic at location left-center.
object 1: camera of color black, texture smooth at location center-right. …’’
2D Representations Refinement   The concise descriptions of identified objects are used as input text prompts for a language-guided segmentation model, enabling the acquisition of $K$ segmentation masks $\{M_{k}^{occ}\}_{k=1}^{K}$, with each mask corresponding to a unique object.
However, an object $i\in[1,K]$ may be occluded by other object(s), leading to an incomplete mask $M_{i}^{occ}$, which may be burdensome when we elevate the image to a 3D representation in the later stage. To resolve this, we create an inpainting mask, $M_{i}^{inp}$, for each object, in which all objects except the one itself are removed and replaced with white pixels. The inpainted masks are again fed to the language-guided segmentation model along with input text prompts such that occlusion-free object masks, $\{M_{k}^{of}\}_{k=1}^{K}$, are obtained. This two-step segmentation process for object $i$ is formulated as:
where seg denotes the language-guided segmentation and $\tau_{i}$ denotes the description of object $i$. In practice, to cleanly remove objects without residual fragments for inpainting, we apply dilation to and expand the white areas. Inpainted background $I_{bg}$ is acquired by removing and replacing all objects with dilated white pixels.
3.2 Coarse 3D Scene Understanding
Scene Size Estimation   Using the estimated metric depth [29] and estimated camera intrinsic matrix by finding field of view (FOV) through perspective fields [29], we backproject to determine the dimensions of the 3D spatial scene.
Prompting: Perspective Canonicalization   3D information without any knowledge regarding the camera perspectives lead to ambiguities [12]. Consider a question “What is the orientation of the bowl relative to the laptop?” with the input scene in Fig. 2, but taken from a top-down perspective. VLMs may output ‘‘downward to the left’’, but the correct answer should be "front-left" because humans perceive orientation from a horizontal angle. To address this, we provide the VLM with $I$, estimated scene size, and maximum and minimum dimensions, allowing it to reason about the camera shot angle (horizontal/top-down/bottom-up). Scene size information helps differentiate shot angles by providing clues about spatial layout and object proportions. For instance, if the depth variation is small, the VLM can infer a top-down or bottom-up angle along with visual cues.